OpenStudy (anonymous):

Solve 7 1/9 - 4 2/3 Show your work.

OpenStudy (anonymous):

64/9 - 14/3 = 64/9 - 42/9 = 22/9 = 2 4/9

OpenStudy (amistre64):

7 + 1/9 - 4 - 2/3 = N 21 + 1/3 - 12 - 2 = 3N 63 + 1 - 36 - 6 = 6N 64 - 42 = 6N 22 = 6N 22/6

OpenStudy (amistre64):

almost had it lol; 3*3 = 9 not 6 ... 22/9

OpenStudy (anonymous):

that's a novel way of dealing with fractions - i like it!

OpenStudy (amistre64):

its a bit much for simpler stuff; but comes in handy when you dont want to fight for a least common denominator :)

OpenStudy (amistre64):

... and dont have access to a calculator lol

OpenStudy (anonymous):

22/9 is okay.